1) use four fours to create 65, and you can use any mathematical symbols, but you can only use four fours as your digits.
2) find four #'s, the sum of which is 45, so that if 2 is added to the first #, 2 is subtracted from the second #, the third is multiplied by 2 and the fourth # is divided by 2, then thr four #'s produced are all the same.

Answer by AnlytcPhil(1806)

1) use four fours to create 65, and you can use any mathematical
symbols, but you can only use four fours as your digits.
`
(44+4)/4 = 65
`
2) find four #'s, the sum of which is 45, so that if 2 is added to
the first #, 2 is subtracted from the second #, the third is
multiplied by 2 and the fourth # is divided by 2, then thr four
#'s produced are all the same.
`
Let the numbers be w, x, y, and z
`
w + x + y + z = 45
w+2 = x-2 = 2y = w/2
`
We can get a system of 4 equations in 4 unknowns:
`
w + x + y + z = 45
w + 2 = x - 2
x - 2 = 2y
2y = w/2
`
Simplify the second one, clear the last one of fractions
by multiplying thru by 2
`
w + x + y + z = 45
w + 4 = x
x - 2 = 2y
4y = w
`
which can be written:
`
`w + x + y + z = 45
`w - x ` ` ` ` = -4
`` ` x - 2y` ` =` 2
-w` ` `+ 4y `` =` 0
`
Can you solve that system? If not ask again.
`
w = -4, x = 0, y = -1, z = 50
`
Edwin